OpenStudy (anonymous):

the directions are" Solve" .... lnx-ln(3x-5)=2

OpenStudy (anonymous):

x/(3x-5)=ln 2

OpenStudy (amistre64):

ln(x/(3x-5)) = 2 x/(3x-5) = e^2

OpenStudy (watchmath):

\(\ln\frac{x}{3x-5}=\ln(e^2)\) \(\frac{x}{3x-5}=e^2\) \(x=e^2(3x-5)\) \((1-3e^2)x=-5e^2\) \(x=\frac{5e^2}{3e^2-1}\)

OpenStudy (anonymous):

Thanks everyone! :D